### v3

# When Parts Are Greater Than Sums: Individual LLM Components Can Outperform Full Models

Abstract
This paper studies in-context learning by decomposing the output of large language models into the individual contributions of attention heads and MLPs (components). We observe curious components: good-performing ones that individually do well on a classification task, even when the full model performs poorly; bad-performing ones that do much worse than chance; and label-biased components that always predict the same label. We find that component accuracies are well-correlated across different demonstration sets and perturbations of prompt templates. Based on our findings, we propose component reweighting, which learns to linearly re-scale the component activations from a few labeled examples. Given $24$ labeled examples, our method improves by an average of $6.0\%$ accuracy points over $24$-shot ICL across 8 tasks on Llama-2-7B. Overall, this paper both enriches our understanding of ICL and provides a practical method for improvement by examining model internals.
When Parts Are Greater Than Sums:
Individual LLM Components Can Outperform Full Models
Ting-Yun Chang   Jesse Thomason   Robin Jia University of Southern California, Los Angeles, CA, USA {tingyun, jessetho, robinjia}@usc.edu
1 Introduction
The rapid progress in large language models (LLMs) has popularized prompting, which guides LLMs to perform tasks with instructions or examples. Notably, in-context learning (ICL; Brown et al., 2020) adapts LLMs to a new task using only a few labeled examples without parameter updates. However, how LLMs react to the in-context examples is sometimes unintuitive Min et al. (2022b). Recently, Sclar et al. (2024) and Voronov et al. (2024) find that even for instruction-tuned (Ouyang et al., 2022) or very large models, adding a space or newline in prompts can greatly affect accuracy.
Figure 1: Each dot represents a component (attention head or MLP) under 4-shot ICL on Llama-2-7B. The $x$-axis shows how often a component predicts “positive” on the test set. Up: We discover good-performing (blue), bad-performing (red), and label-biased (green) components. Down: Most components identified on SST2 show similar characteristics on Yelp-polarity.
We look into the LLM internals to understand what causes the surprising behavior across various ICL settings. Our work stands in contrast to prior studies, which often treat LLMs as black boxes and alter either the input Chen et al. (2023); Bertsch et al. (2024) or output Zhao et al. (2021); Holtzman et al. (2021). We introduce a new view of ICL by decomposing the output of an LLM into the sum of individual contributions of MLPs and attention heads, denoted “components.” Figure 1 reveals three types of curious components: good-performing ones (blue) that individually perform well or even outperform the full model, bad-performing ones (red) that perform below chance, and label-biased ones (green) that predict the same label on the entire test set. We observe these three classes of components on Llama-2-7B, Llama-2-13B Touvron et al. (2023), Llama-3-8B Dubey et al. (2024), and Mistral-Instruct-7B Jiang et al. (2023) across 8 classification tasks.
We study the sensitivity of LLM components to multiple prompts formed by different demonstrations and templates. We also construct contrast sets of templates—pairs of similar templates that yield large differences in ICL accuracy. Despite large variance in full-model accuracy, we find that component accuracies correlate well across different demonstrations ($r=0.80$ on average) and contrast set templates ($r=0.57$). The top-performing components in contrast set pairs overlap and achieve decent accuracy even when the full model performs near random (Figure 2). Nonetheless, the component accuracies of two sampled templates are less correlated ($r=0.34$). Further, good-performing components generalize well to out-of-distribution test sets. For instance, the top-1 component for MNLI outperforms the full Llama-2-13B model by $9.1\%$ on MedNLI; Figure 1 also shows that components are transferrable from SST2 to Yelp. We conclude that components are relatively consistent in their behavior across prompts and datasets.
Inspired by our findings, we propose component reweighting. Compared to prior work that selects prompts from a large pool of labeled data to improve ICL accuracy Liu et al. (2022b), component reweighting softly selects components by learning weights from few-shot examples to scale component activations. Training these weights only involves learning a linear layer, which takes less than a minute on one CPU. Overall, component reweighting better utilizes the same labeled examples, improving over $24$-shot ICL by $6.0\%,2.2\%,5.1\%,1.6\%$ on Llama-2-7B, Llama-2-13B, Mistral-Instruct-7B, and Llama-3-8B, respectively. At the same time, it enjoys similar inference speed as 4-shot ICL.
Finally, we study the training dynamics of components using the Pythia pretraining checkpoints Biderman et al. (2023). During pretraining, good-performing components emerge well before the full model performs well. These findings suggest that LLMs acquire the internal ability to perform ICL early in training, but this ability only surfaces in the full model’s behavior later on.
Overall, our work conducts extensive analysis of LLM internals, which motivates a practical method to improve ICL. We hope to inspire future work that further sheds light on LLM internals in order to improve performance. Our implementation is available at https://github.com/terarachang/LLMDecomp.
2 Decomposing the Transformer in ICL
We introduce a new view of in-context learning by decomposing the Transformer architecture Vaswani et al. (2017). Our decomposition is exact—a mathematically equivalent formula for the model’s outputs—and enables us to analyze model internals without training additional parameters (unlike, e.g., probing). We first discuss what our new view offers over the standard view of ICL, and then walk through the mathematical details.
2.1 A New View of In-Context Learning
Standard view.
An LLM performs in-context learning (ICL) on a task based on a few demonstrations without training, where each demonstration is a templated example $(x,y)$ consisting of an input $x$ and a label word $y$. We refer to a sequence of $K$ demonstrations $[x_{1},y_{1},\dotsc,x_{K},y_{K}]$ as a prompt. The LLM makes predictions on a test input $x_{\text{test}}$ conditioned on the prompt, denoted by ${\arg\max}_{y\in\mathcal{Y}}P(y|\emph{prompt},x_{\text{test}})$, where $\mathcal{Y}$ is the set of possible label words in a classification task.
Our view.
The residual stream of an LLM directly carries the information of the initial hidden state, every attention head, and every MLP, collectively named “components,” towards the output layer. We view this information as the direct contributions [^1] In comparison, a component has indirect contributions to the output by affecting other components in later layers Wang et al. (2023a). This paper focuses on direction contributions. of components to the output logits, and derive a formula for logits, $\sum_{j}\mathbf{g}_{j}$, where $\mathbf{g}_{j}$ is the direct contribution of the component indexed by $j$. We can obtain the predictions of component $j$ with ${\arg\max}_{y\in\mathcal{Y}}\;\mathbf{g}_{j}$, and then calculate its individual ICL accuracy. Specifically, we derive $\mathbf{g}_{j}=U\cdot C_{j}$ in Eq. 8 below, where $U$ is the output embedding matrix and $C_{j}$ is the post-layernorm activations of component $j$. We name the operation $(C_{j}\mapsto U\cdot C_{j})$ as early decode, sharing the same spirit as nostalgebraist (2020) and Geva et al. (2022), which interpret hidden representations by decoding through $U$. Compared to the standard view, we can directly study the behavior of individual components (Figure 2), characterizing them and scaling their contributions to the model output.
Figure 2: Left: Transformer decomposition. The components—MLPs and attention heads—are filled with blue, and the blue lines show the flow of early decoding. Right: We can calculate the individual accuracy of every component after decomposition. Although a pair of templates that only differ slightly yield very different accuracies ($0.39$ vs. $0.89$ on AGNews with Llama-2-7B), the accuracies of their internal components are highly correlated. The top components for Template 1 overlap with the ones for Template 2 and achieve $>0.7$ accuracy despite the poor full-model accuracy.
2.2 A Walkthrough of the Decomposition
A Transformer of $L$ layers consists of a multi-headed attention (MHA) and MLP in every layer. Let $a^{(l)}\in\mathbb{R}^{d}$ and $m^{(l)}\in\mathbb{R}^{d}$ be the output of the MHA and MLP at layer $l$, respectively. Due to residual connections, the hidden state $x^{(l)}\in\mathbb{R}^{d}$ is:
$x^{(l)}$ $=x^{(l-1)}+a^{(l)}+m^{(l)},$ (1)
$x^{(L)}$ $=x^{(0)}+\sum_{l=1}^{L}\;\left(a^{(l)}+m^{(l)}\right).$ (2)
Note that GPT2-like LLMs apply layernorm before MHA and MLP Radford et al. (2019); thus, layernorm is already taken into account as part of the formula for computing $a^{(l)}$ and $m^{(l)}$ (see A.3).
An MHA $a^{(l)}$ is composed of $n$ attention heads:
$a^{(l)}=W_{o}^{(l)}\cdot\text{Concat}([h^{(l)}_{1},\dots,h^{(l)}_{n}])$ (3)
for $h^{(l)}_{i}\in\mathbb{R}^{d_{\text{head}}}$ a head and $W_{o}^{(l)}\in\mathbb{R}^{d\times nd_{\text{head}}}$ the output projection in MHA aggregating all heads. Elhage et al. (2021) rewrite Eq. 3 by segmenting $W_{o}^{(l)}$ into $n$ matrices $W_{o_{i}}^{(l)}\in\mathbb{R}^{d\times d_{\text{head}}}$:
$a^{(l)}$ $=\sum_{i=1}^{n}\left(W_{o_{i}}^{(l)}\cdot h^{(l)}_{i}\right)=\sum_{i=1}^{n}\tilde{h}_{i}^{(l)},$ (4)
$\text{where}\;[W_{o_{1}}^{(l)},\dots,W_{o_{n}}^{(l)}]=W_{o}^{(l)}$ (5)
Thus, we can treat each head as a single component adding $\tilde{h}_{i}^{(l)}=W_{o_{i}}^{(l)}\cdot h^{(l)}_{i}$ to the residual stream.
SST2 BoolQ QQP WiC RTE MNLI AGNews ARC-Easy Avg.
Llama2 7B Full $75.8_{\hskip 1.42271pt18.1}$ $69.2_{\hskip 1.42271pt12.0}$ $61.3_{\hskip 1.42271pt9.9}$ $52.4_{\hskip 1.42271pt3.0}$ $\textbf{68.9}_{\hskip 1.42271pt3.2}$ $34.4_{\hskip 1.42271pt1.7}$ $70.0_{\hskip 1.42271pt19.9}$ $\textbf{57.5}_{\hskip 1.42271pt14.4}$ $61.2$
Oracle-T1 $\textbf{91.7}_{\hskip 1.42271pt\phantom{1}0.9}$ $\textbf{69.7}_{\hskip 1.42271pt\phantom{1}7.7}$ $\textbf{67.8}_{\hskip 1.42271pt4.3}$ $\textbf{57.8}_{\hskip 1.42271pt1.1}$ $64.6_{\hskip 1.42271pt2.7}$ $\textbf{46.3}_{\hskip 1.42271pt3.3}$ $\textbf{80.8}_{\hskip 1.42271pt\phantom{1}5.2}$ $54.5_{\hskip 1.42271pt10.1}$ 66.6
Oracle-B1 $12.1_{\hskip 1.42271pt\phantom{1}2.7}$ $34.1_{\hskip 1.42271pt\phantom{1}7.3}$ $32.5_{\hskip 1.42271pt3.9}$ $42.9_{\hskip 1.42271pt1.2}$ $34.7_{\hskip 1.42271pt2.8}$ $24.1_{\hskip 1.42271pt2.4}$ $\phantom{1}3.0_{\hskip 1.42271pt\phantom{1}1.1}$ $12.7_{\hskip 1.42271pt\phantom{1}4.2}$ $24.5$
Llama2 13B Full $89.0_{\hskip 1.42271pt5.3}$ $\textbf{77.6}_{\hskip 1.42271pt6.8}$ $71.0_{\hskip 1.42271pt6.8}$ $55.0_{\hskip 1.42271pt3.8}$ $75.1_{\hskip 1.42271pt2.3}$ $45.7_{\hskip 1.42271pt7.9}$ $70.8_{\hskip 1.42271pt20.6}$ $\textbf{73.2}_{\hskip 1.42271pt13.7}$ $69.7$
Oracle-T1 $\textbf{92.5}_{\hskip 1.42271pt0.6}$ $77.5_{\hskip 1.42271pt6.0}$ $\textbf{73.5}_{\hskip 1.42271pt2.9}$ $\textbf{60.4}_{\hskip 1.42271pt1.2}$ $\textbf{75.7}_{\hskip 1.42271pt2.3}$ $\textbf{56.4}_{\hskip 1.42271pt4.7}$ $\textbf{84.6}_{\hskip 1.42271pt\phantom{1}3.6}$ $73.1_{\hskip 1.42271pt\phantom{1}7.9}$ 74.2
Oracle-B1 $\phantom{1}8.2_{\hskip 1.42271pt1.0}$ $27.1_{\hskip 1.42271pt9.7}$ $31.8_{\hskip 1.42271pt3.4}$ $39.5_{\hskip 1.42271pt1.6}$ $27.9_{\hskip 1.42271pt2.8}$ $18.6_{\hskip 1.42271pt2.6}$ $\phantom{1}1.8_{\hskip 1.42271pt\phantom{1}0.9}$ $\phantom{1}5.4_{\hskip 1.42271pt\phantom{1}3.5}$ $20.0$
Mistral Ins 7B Full $90.1_{\hskip 1.42271pt2.9}$ $\textbf{81.3}_{\hskip 1.42271pt2.1}$ $70.9_{\hskip 1.42271pt7.2}$ $58.5_{\hskip 1.42271pt4.2}$ $80.5_{\hskip 1.42271pt1.7}$ $56.1_{\hskip 1.42271pt5.0}$ $83.0_{\hskip 1.42271pt5.7}$ $\textbf{79.8}_{\hskip 1.42271pt1.4}$ $75.0$
Oracle-T1 $\textbf{91.9}_{\hskip 1.42271pt0.7}$ $80.8_{\hskip 1.42271pt2.0}$ $\textbf{75.6}_{\hskip 1.42271pt2.6}$ $\textbf{60.6}_{\hskip 1.42271pt2.2}$ $\textbf{81.3}_{\hskip 1.42271pt0.8}$ $\textbf{61.5}_{\hskip 1.42271pt3.3}$ $\textbf{83.7}_{\hskip 1.42271pt4.3}$ $78.5_{\hskip 1.42271pt2.2}$ 76.7
Oracle-B1 $\phantom{1}8.1_{\hskip 1.42271pt0.9}$ $19.5_{\hskip 1.42271pt2.5}$ $25.8_{\hskip 1.42271pt4.1}$ $39.3_{\hskip 1.42271pt2.8}$ $20.0_{\hskip 1.42271pt1.7}$ $14.6_{\hskip 1.42271pt2.9}$ $\phantom{1}1.8_{\hskip 1.42271pt0.7}$ $\phantom{1}4.6_{\hskip 1.42271pt1.3}$ $16.7$
Llama3 8B Full $91.4_{\hskip 1.42271pt1.7}$ $\textbf{79.2}_{\hskip 1.42271pt7.2}$ $74.0_{\hskip 1.42271pt8.0}$ $58.7_{\hskip 1.42271pt4.7}$ $\textbf{76.5}_{\hskip 1.42271pt2.2}$ $59.4_{\hskip 1.42271pt3.7}$ $\textbf{84.0}_{\hskip 1.42271pt6.6}$ $\textbf{87.4}_{\hskip 1.42271pt5.5}$ $76.3$
Oracle-T1 $\textbf{92.3}_{\hskip 1.42271pt1.0}$ $77.4_{\hskip 1.42271pt7.3}$ $\textbf{77.4}_{\hskip 1.42271pt3.7}$ $\textbf{64.5}_{\hskip 1.42271pt2.7}$ $76.3_{\hskip 1.42271pt2.9}$ $\textbf{60.7}_{\hskip 1.42271pt1.5}$ $81.4_{\hskip 1.42271pt5.7}$ $86.0_{\hskip 1.42271pt5.9}$ 77.0
Oracle-B1 $\phantom{1}9.0_{\hskip 1.42271pt0.9}$ $22.5_{\hskip 1.42271pt7.4}$ $23.5_{\hskip 1.42271pt3.9}$ $36.7_{\hskip 1.42271pt3.3}$ $23.4_{\hskip 1.42271pt2.1}$ $10.0_{\hskip 1.42271pt4.2}$ $\phantom{1}1.4_{\hskip 1.42271pt0.6}$ $\phantom{1}1.9_{\hskip 1.42271pt0.9}$ $16.0$
Random $50.0$ $50.0$ $50.0$ $50.0$ $50.0$ $33.3$ $25.0$ $25.0$ $41.7$
Table 1: $\{3,4\}$-shot ICL accuracy of 8 tasks and the average accuracy (Avg.). We run 15 prompts for each task (see section 3) and report the mean accuracy and standard deviation. We show the existence of good components (Oracle-T1) inside LLMs that individually perform on par with the full model (Full) on diverse tasks. Similarly, there exist bad components (Oracle-B1) that perform substantially below chance (Random).
Finally, through the output embedding matrix $U\in\mathbb{R}^{|\text{Vocab}|\times d}$, the output logits are:
$\text{logits}=U\cdot\text{LN}(x^{(L)})$
$=U\cdot\text{LN}\left(x^{(0)}+\sum_{l=1}^{L}\sum_{i=1}^{n}\tilde{h}_{i}^{(l)}+\sum_{l=1}^{L}m^{(l)}\right)$
$=U\cdot\text{LN}\left(\sum_{j=1}^{1+L\times n+L}z_{j}\right),$ (6)
where $z=[x^{(0)},\tilde{h}_{1}^{(1)},\dotsc,\tilde{h}_{n}^{(L)},m^{(1)},\dotsc,m^{(L)}]$ in Eq. 6 and we index every term in the summation with $j$. LN()($\cdot$)( ⋅ ) denotes the final layernorm, specifically, RMSNorm Zhang and Sennrich (2019) for LLMs in our paper (see A.3). In Eq. 6, LN$(\sum_{j}z_{j})=\frac{\sum_{j}z_{j}}{\text{RMS}(\sum_{j}z_{j})}\odot\gamma$, where RMS denotes root mean square, $\odot$ denotes element-wise multiplication, and $\gamma\in\mathbb{R}^{d}$ is the affine parameters. By pre-computing $\hat{\gamma}=\frac{\gamma}{\text{RMS}(\sum_{j}z_{j})}$, we have:
$\text{logits}=U\cdot\left(\sum_{j}z_{j}\odot\hat{\gamma}\right)$ (7)
$=\sum_{j}U\cdot C_{j},\;\text{where}\;C_{j}=z_{j}\odot\hat{\gamma}$ (8)
We refer to all $C_{j}\in\mathbb{R}^{d}$ as the component activations, which include the activations of attention heads and MLPs after the final layernorm. [^2] Empirically, we find that $x^{(0)}$ has near-random ICL accuracy on all the tasks, so we omit it in the rest of the paper. Now that we have broken down the Transformer output into simple additions in Eq. 8, we can easily analyze the direct contribution of each component to the logits through the residual stream, $\mathbf{g}_{j}=U\cdot C_{j}$.
In ICL, we only need to do the decomposition when LLMs start to generate, i.e, when processing the last token of the input. The computations on the other tokens are the same as the standard ICL. In all our experiments, we use single-token label words. We use multiple templates from Bach et al. (2022) that cover diverse label words for each task.
3 Characterizing Components for ICL
We conduct in-context learning across 8 classification tasks on 4 LLMs: Llama-2-7B, Llama-2-13B, Mistral-Instruct-7B, and Llama-3-8B. ICL is sensitive to prompts, so we randomly sample 5 disjoint sets of demonstrations formatted with 3 templates and report the standard deviation across the 15 runs. To avoid majority and recency biases Zhao et al. (2021), each prompt consists of the same number of demonstrations from every class in shuffled order. We use $K=3$ demonstrations for 3-way classification tasks and $K=4$ for the other tasks. Except for section 5.1, we refer to $K=\{3,4\}$ without further notice. We sample 2000 examples with balanced labels as the test set for every task. Please see A.1 for details about the tasks and templates.
3.1 Good and Bad-Performing Components
Across all the tasks and LLMs, we observe good-performing components that perform well or even outperform the full model, and bad-performing components that individually perform much worse than chance (blue and red dots in Figure 1, respectively). Table 1 compares the full model (Full) with the top-1 (Oracle-T1) and bottom-1 (Oracle-B1) components selected on the test set. On average, Oracle-T1 outperforms Full by $5.4\%,4.5\%,1.7\%,0.7\%$ on Llama-2-7B, Llama-2-13B, Mistral-Instruct-7B, and Llama-3-8B, respectively; Oracle-B1 underperforms random guessing (Random) by $17.2\%,20.7\%,25.0\%,25.7\%$.
3.2 Label-Biased Components
Besides good and bad-performing components, we also observe label-biased components, which predict a certain label on the entire test set (the green dots in Figure 1). These components exist in all the tasks and LLMs we study, accounting for $29.1\%,26.4\%,22.8\%,29.7\%$ of components on average in Llama-2-7B, Llama-2-13B, Mistral-Instruct-7B, and Llama-3-8B, respectively (Table 5). In A.2, we show that even when we prompt the model with all demonstrations of positive labels, the most biased component still insists on predicting “negative” on the entire test set, and vice versa.
3.3 Mechanistic Understanding of Bad-Performing Heads
Prior work studies the mechanism of certain components in LLMs, showing that there are negative mover attention heads that write in the opposite direction of the expected answer Wang et al. (2023a) and copy suppression heads that suppress the prediction of a prior token in the context McDougall et al. (2023). Inspired by them, we investigate the mechanism behind bad-performing heads identified by our decomposition. We focus on label tokens in the context, as Wang et al. (2023b) show that label words serve as anchors in ICL.
We conduct a case study on Llama-2-7B with 4-shot balanced in-context examples from SST2. We examine the bottom-5 attention heads that have the worst ICL accuracy on SST2. We find that three of these heads, L19H15, L15H14, and L18H9, assign top attention probabilities to all 4 label tokens of the 4-shot in-context examples when predicting test examples. Furthermore, despite their poor ICL accuracy, these heads actually assign higher attention to the correct in-context label tokens than the incorrect ones most of the time ($>70\%$ of the test examples). In other words, when a test example has a positive label, these heads assign higher attention [^3] We average the attention probabilities of the same label tokens and then compare the average ones of the two labels. to the tokens “positive” in the context than the tokens “negative”. We also observe that the more the heads attend to “positive” in the context, the lower the inner product between the head and the output embedding of the token “positive”, with the correlation $r=-0.97,-0.96,-0.89$ for L19H15, L15H14, and L18H9, respectively.
[…]
SST2 BoolQ QQP AGNews ARC
Corr (1) Demo 0.81 0.84 0.60 0.89 0.88
(2) Temp 0.40 0.16 0.03 0.68 0.44
(3) Cst T 0.72 0.63 0.23 0.82 0.46
IoU (1) Demo 0.36 0.74 0.27 0.63 0.70
(2) Temp 0.12 0.01 0.01 0.20 0.20
(3) Cst T 0.40 0.23 0.02 0.36 0.45
Table 2: The average correlation and IoU between (1) two random sets of demonstrations, (2) two random templates, and (3) two minimally contrastive templates.
4 Transferability of Components
We observe moderate to high component transferability across demonstrations, minimally contrastive templates, and data distributions, whereas there is little transferability across randomly sampled templates. Our decomposition uncovers hidden abilities of individual components when the full model performs poorly.
4.1 Transfer across Prompt Variants
We first measure the agreement in component accuracies between (1) two disjoint sets of demonstrations with a fixed template, (2) two randomly sampled templates with fixed demonstrations, and (3) two minimally-contrastive templates with fixed demonstrations. Recall that we have 5 sets of demonstrations and 3 templates in total (section 3); here, we calculate the average agreement between every pair. For (3), we construct contrast sets Gardner et al. (2020) by minimally editing the worst-performing template out of the 3 templates into a good template, which yields at least $10\%$ improvement in average accuracy. Our edits include adding a space, removing a newline, or changing label words (see Table 10). We use two metrics to measure the agreement between each pair: Pearson correlation of the accuracies of all components and the intersection over union (IoU) on the sets of top-5 components, which measures whether the top-performing components of the pair overlap.
Table 2 summarizes the results on Llama-2-7B; A.6 shows similar findings on other models. (1) The accuracies of the internal components are highly consistent across different choices of demonstrations, having strong correlations and an average of $0.54$ IoU. (2) The components have much weaker agreement across randomly sampled templates, having a near 0 IoU on BoolQ and QQP. (3) Nevertheless, there is agreement between minimally contrastive templates (Cst T), with an average correlation of $0.57$ across tasks, despite contrasting full-model accuracy. For example, Figure 2 demonstrates that full-model accuracy changes dramatically ($39\%$ vs $89\%$) in a minimal pair of templates, but internal components have a high correlation of $0.81$ and the pair shares top-performing components. Combining (2) and (3) suggests components behave similarly on similar templates, but this similarity decreases as the templates diverge.
4.2 Transfer to Out-of-Distribution Test Sets
We further study whether the best component selected on the test set can still perform well on an out-of-distribution (OOD) test set. We name this method, which uses a single component to make predictions, as Transfer-1. Specifically, we study component transferability from SST2 to Yelp-polarity, MNLI to MedNLI, and BoolQ to BoolQ Contrast Set. We compare Transfer-1 with using the full model (Full) on the OOD test sets. To understand the best possible Transfer-1 accuracy, we also report the best component accuracy directly selected on the OOD set, Oracle-1.
Table 3 shows that Transfer-1 closely matches Oracle-1 overall, suggesting that the top-performing components are transferable across data distribution. Moreover, Transfer-1 sometimes outperforms Full, especially on Llama2 models, showing the hidden abilities of the internal components.
4.3 Transfer between Two Opposite Tasks
We conduct a case study of component transferability across instructions using Task069 and Task070 of Super-NaturalInstructions Wang et al. (2022b), both of which are binary abductive NLI tasks Bhagavatula et al. (2020). The instruction for Task069 asks for correct answers, while Task070 asks for incorrect ones (“pick the one that makes less sense;” see Figure 7 for the full instructions). Examples in the two tasks are not parallel.
We find that Mistral-Instruct-7B achieves good accuracy across 15 runs on Task069 ($76.8\pm 2.4$), but below chance on Task070 ($40.6\pm 5.4$). We observe a strong negative correlation, $r=-0.60$ on average, between the component accuracies of the two tasks. The worst-performing components in Task069 become the top-performing in Task070 and vice versa. The correlation suggests that the model has the ability to solve Task070, but misunderstands negation. Thus, we apply the Transfer-1 method (section 4.2) but select the worst-performing component from Task069 and then calculate its individual accuracy on Task070. Transfer-1 achieves $58.7\pm 4.8$ accuracy across the 15 runs, an improvement of $\mathbf{18.1\%}$ over the full model. These results suggest that components behave consistently even across tasks with opposite instructions, as the active components in Task069 are also active in Task070.
Yelp-polarity MedNLI BoolQ Cst
Llama2 7B Full $84.7_{\hskip 1.42271pt15.4}$ $34.3_{\hskip 1.42271pt1.7}$ $\textbf{64.9}_{\hskip 1.42271pt9.8}$
Transfer-1 $\textbf{94.9}_{\hskip 1.42271pt\phantom{1}3.1}$ $\textbf{42.6}_{\hskip 1.42271pt4.7}$ $64.3_{\hskip 1.42271pt7.9}$
Oracle-1 $96.9_{\hskip 1.42271pt\phantom{1}0.7}$ $48.8_{\hskip 1.42271pt2.3}$ $66.2_{\hskip 1.42271pt5.7}$
Llama2 13B Full $95.9_{\hskip 1.42271pt1.4}$ $46.8_{\hskip 1.42271pt9.6}$ $72.0_{\hskip 1.42271pt7.6}$
Transfer-1 $\textbf{96.0}_{\hskip 1.42271pt1.8}$ $\textbf{55.9}_{\hskip 1.42271pt4.0}$ $\textbf{72.3}_{\hskip 1.42271pt6.5}$
Oracle-1 $97.1_{\hskip 1.42271pt0.4}$ $57.0_{\hskip 1.42271pt3.7}$ $73.0_{\hskip 1.42271pt6.1}$
Mistral Ins 7B Full $\textbf{97.0}_{\hskip 1.42271pt0.5}$ $57.3_{\hskip 1.42271pt5.7}$ $\textbf{74.6}_{\hskip 1.42271pt3.5}$
Transfer-1 $95.6_{\hskip 1.42271pt1.6}$ $\textbf{61.9}_{\hskip 1.42271pt4.8}$ $73.7_{\hskip 1.42271pt3.7}$
Oracle-1 $97.1_{\hskip 1.42271pt0.4}$ $62.7_{\hskip 1.42271pt4.1}$ $74.5_{\hskip 1.42271pt3.6}$
Llama3 8B Full $\textbf{97.8}_{\hskip 1.42271pt0.4}$ $61.0_{\hskip 1.42271pt2.2}$ $\textbf{77.3}_{\hskip 1.42271pt7.5}$
Transfer-1 $95.9_{\hskip 1.42271pt4.4}$ $\textbf{61.3}_{\hskip 1.42271pt0.8}$ $73.9_{\hskip 1.42271pt8.4}$
Oracle-1 $97.9_{\hskip 1.42271pt0.5}$ $61.6_{\hskip 1.42271pt0.6}$ $74.8_{\hskip 1.42271pt8.9}$
Table 3: The average ICL accuracy and standard deviation on OOD test sets. The components selected on the in-distribution test sets (Transfer-1) can transfer to OOD sets, performing similarly to the oracle components (Oracle-1) directly selected on the OOD sets.
5 Component Reweighting
5.1 Proposed Method
Our findings in section 4 show the promising direction of selecting internal components to improve ICL. Therefore, we propose a method that reweights components by learning a weight $w_{j}\in\mathbb{R}$ on every component activation $C_{j}$. Reweighting is a soft version of selection, which can be learned by gradient descent on very few examples.
Given $K$ labeled examples, instead of using all of them as ICL demonstrations, we divide them into a demonstration set $\mathcal{D}_{\text{demo}}$ and a training set $\mathcal{D}_{\text{train}}$. We first randomly sample $K^{\prime}=\{3,4\}$ examples with balanced labels as demonstrations and use the remaining examples as $\mathcal{D}_{\text{train}}$ to train the component weights. Specifically, we can rewrite Eq. 8 as $\text{logits}=\sum_{j}w_{j}(U\cdot C_{j})$, where $w_{j}=1$ for all $j$. Because of the existence of good and bad-performing components, weighing all components equally may not be optimal. Therefore, we tune the weights $w\in\mathbb{R}^{N}$ of $N$ components on $\mathcal{D}_{\text{train}}$ with cross-entropy loss and $L_{1}$ regularization, while keeping the LLM frozen:
$\mathcal{L}=\sum_{(x,y)\in\mathcal{D}_{\text{train}}}-\log P_{rw}(y|x)+\lambda\|w\|_{1},$ (9)
$P_{rw}(y|x)=\text{softmax}\left(\sum_{j=1}^{N}w_{j}\;(U_{\mathcal{Y}}\cdot C_{j})\right)_{y},$
where $U_{\mathcal{Y}}\in\mathbb{R}^{|\mathcal{Y}|\times d}$ is a submatrix of $U$ that comprises the output embeddings of label words, $P_{rw}$ is the probability distribution of the LLM after reweighting, and $\lambda$ is the hyperparameter of the $L_{1}$ loss to encourage sparsity on the component weights. We obtain the activations $\{C_{j}\}_{j=1}^{N}$ of all components in one $K^{\prime}$-shot forward pass, computed on the prompt derived from $\mathcal{D}_{\text{demo}}$, followed by $x$. Our method scales each component’s direct contributions to the logits ($U_{\mathcal{Y}}\cdot C_{j}\in\mathbb{R}^{|\mathcal{Y}|}$) by $w_{j}$. In practice, we cache these contributions on the entire training set as input features to the linear layer $w$, which allows us to discard the entire LLM while training $w$ (line 9 and 13 in Algorithm 1), saving tremendous training time and GPU memory. The cache only requires $O(|\mathcal{Y}|\times N\times|\mathcal{D}_{\text{train}}|)$ space. At inference time, the overhead of our method over $K^{\prime}$-shot ICL is to early decode $N$ components and apply the learned weights, i.e., $\sum_{j=1}^{N}w_{j}\;(U_{\mathcal{Y}}\cdot C_{j})$. As both $|\mathcal{Y}|$ and $N$ are small ($N<2000$ for all LLMs in this paper), the overhead is negligible compared to the computation of the LLM itself.
Algorithm 1 Component Reweighting
1:Input: $K$ labeled examples, a test set $\mathcal{D}_{\text{test}}$, a set of label words $\mathcal{Y}$, an LLM $\mathcal{M}$, the number of components $N$
2:Output: $\mathcal{Z}$, the predictions of $\mathcal{M}$ on $\mathcal{D}_{\text{test}}$
3:Split $K$ examples into a prompt consists of $K^{\prime}$ demonstrations and a training set $\mathcal{D}_{\text{train}}$ of $K-K^{\prime}$ examples
4:$U_{\mathcal{Y}}\leftarrow$ concatenate the output embeddings of $\mathcal{Y}$ in $\mathcal{M}$
5:Initialize $\mathcal{G}^{\text{train}}\leftarrow\varnothing$
6:for  $(x,y)\in\mathcal{D}_{\text{train}}$  do
7:     $\{C_{j}\}_{j=1}^{N}\leftarrow$ $\mathcal{M}(\emph{prompt},x)$triangleright $K^{\prime}$-shot ICL
8:     for  $j\leftarrow 1$ to $N$  do
9:         $\mathcal{G}^{\text{train}}\leftarrow\mathcal{G}^{\text{train}}\cup(U_{\mathcal{Y}}\cdot C_{j})$triangleright early decode
10:     end for
11:end for
12:Initialize $w\leftarrow[1,\dotsc,1]\in\mathbb{R}^{N}$
13:Train the weights $w$ on $\mathcal{G}^{\text{train}}$ with Eq. 9
14:Initialize $\mathcal{Z}\ \leftarrow\varnothing$triangleright Start Inference
15:for  $(x,y)\in\mathcal{D}_{\text{test}}$  do
16:     $\{C_{j}\}_{j=1}^{N}\leftarrow$ $\mathcal{M}(\emph{prompt},x)$triangleright $K^{\prime}$-shot ICL
17:     Initialize $\mathbf{g}\leftarrow[0,\dotsc,0]\in\mathbb{R}^{|\mathcal{Y}|}$
18:     for  $j\leftarrow 1$ to $N$  do $\triangleright$ Test-Time Overhead
19:         $\mathbf{g}\leftarrow\mathbf{g}+w_{j}(U_{\mathcal{Y}}\cdot C_{j})$triangleright early decode
20:     end for
21:     $\hat{y}\leftarrow{\arg\max}_{y\in\mathcal{Y}}\;\mathbf{g}$
22:     $\mathcal{Z}\ \leftarrow\mathcal{Z}\ \cup\hat{y}$
23:end for
24:return $\mathcal{Z}$
5.2 Baselines
Standard ICL.
The simplest baseline is to use all the $K$ labeled examples as demonstrations. Since the other methods use $K^{\prime}$ examples as demonstrations, we report the accuracy of standard $K^{\prime}$-shot ICL using the same $\mathcal{D}_{\text{demo}}$ for reference.
Prompt Selection.
Liu et al. (2022b) improve ICL accuracy by selecting demonstrations from a pool of labeled data for each test example. Here, we select from the given $K$ labeled examples. Following Rubin et al. (2022), we use SBERT Reimers and Gurevych (2019) to encode examples into sentence embeddings and select the $K^{\prime}=\{3,4\}$ nearest neighbors under cosine similarity as the demonstrations for each test example.
Calibration.
As LLMs tend to predict a certain class over others, Zhao et al. (2021) reweight the output class probabilities. They use context-free inputs, such as “N/A”, to calibrate the probability distribution. However, Fei et al. (2023) and Zhou et al. (2023) find context-free inputs sometimes ineffective, because in-domain context is important for calibration. Thus, we introduce Calib+, which calibrates the original probabilities $\mathbf{p}\in\mathbb{R}^{|\mathcal{Y}|}$ with a training set of in-distribution labeled examples, $\mathcal{D}_{\text{train}}$. We train the calibration weights $\mathbf{v}\in\mathbb{R}^{|\mathcal{Y}|}$ on $\mathcal{D}_{\text{train}}$ with cross-entropy loss and obtain the calibrated probabilities $\hat{\mathbf{p}}=\text{softmax}(\mathbf{v}\cdot\mathbf{p})$. For direct comparisons, we split the $K$ examples into the same $\mathcal{D}_{\text{demo}}$ and $\mathcal{D}_{\text{train}}$ sets as component reweighting for Calib+, where $|\mathcal{D}_{\text{demo}}|=K^{\prime}$. We include the training details of both methods in A.8.
SST2 BoolQ QQP WiC RTE MNLI AGNews ARC-Easy Avg.
Llama-2-7B Standard $3,4$ $75.8_{\hskip 1.42271pt18.1}$ $69.2_{\hskip 1.42271pt12.0}$ $61.3_{\hskip 1.42271pt9.9}$ $52.4_{\hskip 1.42271pt3.0}$ $68.9_{\hskip 1.42271pt3.2}$ $34.4_{\hskip 1.42271pt1.7}$ $70.0_{\hskip 1.42271pt19.9}$ $57.5_{\hskip 1.42271pt14.4}$ $61.2$
Standard $12$ $77.8_{\hskip 1.42271pt19.6}$ $\textbf{71.6}_{\hskip 1.42271pt\phantom{1}8.0}$ $63.6_{\hskip 1.42271pt7.8}$ $52.5_{\hskip 1.42271pt2.4}$ $\textbf{71.1}_{\hskip 1.42271pt2.1}$ $37.0_{\hskip 1.42271pt2.8}$ $69.0_{\hskip 1.42271pt20.8}$ $\textbf{59.6}_{\hskip 1.42271pt13.9}$ $62.8$
PromptS $12$ $73.8_{\hskip 1.42271pt19.2}$ $69.4_{\hskip 1.42271pt10.5}$ $62.2_{\hskip 1.42271pt6.1}$ $53.1_{\hskip 1.42271pt2.7}$ $65.5_{\hskip 1.42271pt1.8}$ $35.5_{\hskip 1.42271pt1.6}$ $59.1_{\hskip 1.42271pt28.7}$ $58.7_{\hskip 1.42271pt11.9}$ $59.7$
Calib12121212 $85.1_{\hskip 1.42271pt\phantom{1}6.0}$ $69.2_{\hskip 1.42271pt13.6}$ $\textbf{73.6}_{\hskip 1.42271pt6.1}$ $55.1_{\hskip 1.42271pt5.1}$ $70.3_{\hskip 1.42271pt2.7}$ $45.5_{\hskip 1.42271pt7.8}$ $77.8_{\hskip 1.42271pt12.2}$ $58.6_{\hskip 1.42271pt14.6}$ $66.9$
CompRW $12$ $\textbf{88.5}_{\hskip 1.42271pt\phantom{1}2.8}$ $70.4_{\hskip 1.42271pt11.2}$ $71.4_{\hskip 1.42271pt5.4}$ $\textbf{56.3}_{\hskip 1.42271pt3.4}$ $70.0_{\hskip 1.42271pt2.8}$ $\textbf{48.3}_{\hskip 1.42271pt4.8}$ $\textbf{87.4}_{\hskip 1.42271pt\phantom{1}2.3}$ $58.3_{\hskip 1.42271pt13.6}$ 68.8
Standard $24$ $77.8_{\hskip 1.42271pt19.5}$ $71.6_{\hskip 1.42271pt\phantom{1}7.3}$ $66.4_{\hskip 1.42271pt5.0}$ $53.2_{\hskip 1.42271pt3.3}$ $\textbf{71.9}_{\hskip 1.42271pt1.5}$ $39.9_{\hskip 1.42271pt3.6}$ $71.1_{\hskip 1.42271pt20.0}$ $58.3_{\hskip 1.42271pt16.2}$ $63.8$
PromptS $24$ $74.2_{\hskip 1.42271pt20.4}$ $68.9_{\hskip 1.42271pt10.2}$ $62.1_{\hskip 1.42271pt4.9}$ $53.6_{\hskip 1.42271pt1.9}$ $64.8_{\hskip 1.42271pt0.9}$ $36.4_{\hskip 1.42271pt1.5}$ $57.5_{\hskip 1.42271pt30.2}$ $58.0_{\hskip 1.42271pt12.5}$ $59.4$
Calib24242424 $87.6_{\hskip 1.42271pt\phantom{1}5.0}$ $70.3_{\hskip 1.42271pt11.9}$ $\textbf{73.4}_{\hskip 1.42271pt5.5}$ $55.8_{\hskip 1.42271pt4.9}$ $70.4_{\hskip 1.42271pt2.7}$ $46.4_{\hskip 1.42271pt6.7}$ $78.4_{\hskip 1.42271pt11.8}$ $\textbf{59.2}_{\hskip 1.42271pt14.4}$ $67.7$
CompRW $24$ $\textbf{90.6}_{\hskip 1.42271pt\phantom{1}1.7}$ $\textbf{71.7}_{\hskip 1.42271pt\phantom{1}9.4}$ $71.9_{\hskip 1.42271pt4.4}$ $\textbf{57.1}_{\hskip 1.42271pt3.0}$ $70.0_{\hskip 1.42271pt4.1}$ $\textbf{49.8}_{\hskip 1.42271pt4.0}$ $\textbf{88.1}_{\hskip 1.42271pt\phantom{1}2.1}$ $58.8_{\hskip 1.42271pt13.6}$ 69.8
Llama-2-13B Standard $3,4$ $89.0_{\hskip 1.42271pt\phantom{1}5.3}$ $77.6_{\hskip 1.42271pt6.8}$ $71.0_{\hskip 1.42271pt6.8}$ $55.0_{\hskip 1.42271pt3.8}$ $75.1_{\hskip 1.42271pt2.3}$ $45.7_{\hskip 1.42271pt7.9}$ $70.8_{\hskip 1.42271pt20.6}$ $73.2_{\hskip 1.42271pt13.7}$ $69.7$
Standard $12$ $\textbf{91.3}_{\hskip 1.42271pt\phantom{1}1.9}$ $78.1_{\hskip 1.42271pt7.4}$ $70.5_{\hskip 1.42271pt7.3}$ $\textbf{59.6}_{\hskip 1.42271pt2.4}$ $74.4_{\hskip 1.42271pt3.5}$ $55.1_{\hskip 1.42271pt6.2}$ $84.7_{\hskip 1.42271pt\phantom{1}7.8}$ $71.2_{\hskip 1.42271pt16.4}$ $73.1$
PromptS $12$ $83.8_{\hskip 1.42271pt10.2}$ $74.9_{\hskip 1.42271pt6.6}$ $64.6_{\hskip 1.42271pt5.7}$ $57.0_{\hskip 1.42271pt2.1}$ $69.5_{\hskip 1.42271pt3.5}$ $48.1_{\hskip 1.42271pt5.4}$ $64.4_{\hskip 1.42271pt29.6}$ $74.2_{\hskip 1.42271pt\phantom{1}9.3}$ $67.1$
Calib12121212 $89.4_{\hskip 1.42271pt\phantom{1}3.2}$ $\textbf{78.4}_{\hskip 1.42271pt6.1}$ $72.1_{\hskip 1.42271pt4.1}$ $58.1_{\hskip 1.42271pt5.1}$ $75.3_{\hskip 1.42271pt1.9}$ $57.3_{\hskip 1.42271pt4.5}$ $81.5_{\hskip 1.42271pt\phantom{1}8.7}$ $74.7_{\hskip 1.42271pt\phantom{1}9.3}$ $73.3$
CompRW $12$ $89.1_{\hskip 1.42271pt\phantom{1}3.2}$ $77.7_{\hskip 1.42271pt6.7}$ $\textbf{72.7}_{\hskip 1.42271pt3.3}$ $58.7_{\hskip 1.42271pt4.0}$ $\textbf{76.2}_{\hskip 1.42271pt2.0}$ $\textbf{60.2}_{\hskip 1.42271pt3.7}$ $\textbf{88.1}_{\hskip 1.42271pt\phantom{1}1.7}$ $\textbf{76.2}_{\hskip 1.42271pt\phantom{1}6.8}$ 74.9
Standard $24$ $\textbf{91.9}_{\hskip 1.42271pt\phantom{1}0.6}$ $77.7_{\hskip 1.42271pt8.2}$ $69.5_{\hskip 1.42271pt8.5}$ $60.6_{\hskip 1.42271pt1.6}$ $74.7_{\hskip 1.42271pt3.3}$ $58.2_{\hskip 1.42271pt7.0}$ $85.8_{\hskip 1.42271pt\phantom{1}4.4}$ $69.1_{\hskip 1.42271pt17.7}$ $73.5$
PromptS $24$ $81.9_{\hskip 1.42271pt13.2}$ $75.1_{\hskip 1.42271pt5.7}$ $64.9_{\hskip 1.42271pt4.8}$ $57.3_{\hskip 1.42271pt1.8}$ $69.5_{\hskip 1.42271pt1.7}$ $49.8_{\hskip 1.42271pt5.1}$ $65.2_{\hskip 1.42271pt28.9}$ $74.2_{\hskip 1.42271pt\phantom{1}9.4}$ $67.2$
Calib24242424 $90.7_{\hskip 1.42271pt\phantom{1}2.1}$ $\textbf{78.6}_{\hskip 1.42271pt6.2}$ $73.1_{\hskip 1.42271pt4.3}$ $\textbf{59.5}_{\hskip 1.42271pt3.2}$ $75.9_{\hskip 1.42271pt1.9}$ $58.4_{\hskip 1.42271pt2.8}$ $82.0_{\hskip 1.42271pt\phantom{1}8.4}$ $75.2_{\hskip 1.42271pt\phantom{1}9.1}$ $74.2$
CompRW $24$ $91.0_{\hskip 1.42271pt\phantom{1}1.8}$ $78.2_{\hskip 1.42271pt6.4}$ $\textbf{74.2}_{\hskip 1.42271pt3.1}$ $58.5_{\hskip 1.42271pt4.1}$ $\textbf{77.1}_{\hskip 1.42271pt1.8}$ $\textbf{62.0}_{\hskip 1.42271pt3.7}$ $\textbf{88.8}_{\hskip 1.42271pt\phantom{1}1.4}$ $\textbf{76.1}_{\hskip 1.42271pt\phantom{1}7.2}$ 75.7
Mistral-Instruct-7B Standard $3,4$ $90.1_{\hskip 1.42271pt2.9}$ $81.3_{\hskip 1.42271pt2.1}$ $70.9_{\hskip 1.42271pt7.2}$ $58.5_{\hskip 1.42271pt4.2}$ $80.5_{\hskip 1.42271pt1.7}$ $56.1_{\hskip 1.42271pt5.0}$ $83.0_{\hskip 1.42271pt5.7}$ $79.8_{\hskip 1.42271pt\phantom{1}1.4}$ $75.0$
Standard $12$ $91.4_{\hskip 1.42271pt0.9}$ $81.2_{\hskip 1.42271pt2.2}$ $67.9_{\hskip 1.42271pt8.7}$ $57.7_{\hskip 1.42271pt2.8}$ $79.1_{\hskip 1.42271pt1.6}$ $57.2_{\hskip 1.42271pt3.6}$ $85.4_{\hskip 1.42271pt3.6}$ $77.7_{\hskip 1.42271pt\phantom{1}5.6}$ $74.7$
PromptS $12$ $90.3_{\hskip 1.42271pt2.5}$ $81.1_{\hskip 1.42271pt1.9}$ $68.7_{\hskip 1.42271pt5.8}$ $57.1_{\hskip 1.42271pt2.7}$ $79.1_{\hskip 1.42271pt1.6}$ $56.7_{\hskip 1.42271pt3.2}$ $84.9_{\hskip 1.42271pt3.0}$ $79.0_{\hskip 1.42271pt\phantom{1}3.0}$ $74.6$
Calib12121212 $\textbf{91.5}_{\hskip 1.42271pt1.6}$ $\textbf{81.3}_{\hskip 1.42271pt1.8}$ $\textbf{75.8}_{\hskip 1.42271pt2.6}$ $58.3_{\hskip 1.42271pt6.6}$ $81.0_{\hskip 1.42271pt1.3}$ $61.9_{\hskip 1.42271pt4.7}$ $85.4_{\hskip 1.42271pt4.0}$ $\textbf{79.6}_{\hskip 1.42271pt\phantom{1}1.6}$ $76.9$
CompRW $12$ $89.9_{\hskip 1.42271pt2.7}$ $80.7_{\hskip 1.42271pt2.7}$ $75.1_{\hskip 1.42271pt2.9}$ $\textbf{60.0}_{\hskip 1.42271pt4.9}$ $\textbf{81.1}_{\hskip 1.42271pt1.3}$ $\textbf{64.7}_{\hskip 1.42271pt4.6}$ $\textbf{87.6}_{\hskip 1.42271pt2.1}$ $79.2_{\hskip 1.42271pt\phantom{1}1.2}$ 77.3
Standard $24$ $91.2_{\hskip 1.42271pt1.0}$ $80.8_{\hskip 1.42271pt2.3}$ $65.3_{\hskip 1.42271pt8.4}$ $57.4_{\hskip 1.42271pt4.0}$ $75.6_{\hskip 1.42271pt1.7}$ $56.6_{\hskip 1.42271pt6.5}$ $85.8_{\hskip 1.42271pt4.3}$ $68.8_{\hskip 1.42271pt16.9}$ $72.7$
PromptS $24$ $90.5_{\hskip 1.42271pt2.6}$ $\textbf{81.3}_{\hskip 1.42271pt2.0}$ $68.9_{\hskip 1.42271pt5.6}$ $57.1_{\hskip 1.42271pt2.1}$ $79.1_{\hskip 1.42271pt1.7}$ $57.4_{\hskip 1.42271pt3.1}$ $86.0_{\hskip 1.42271pt2.1}$ $78.7_{\hskip 1.42271pt\phantom{1}3.3}$ $74.9$
Calib24242424 $\textbf{91.6}_{\hskip 1.42271pt1.5}$ $80.9_{\hskip 1.42271pt2.0}$ $76.1_{\hskip 1.42271pt2.4}$ $59.5_{\hskip 1.42271pt5.4}$ $81.2_{\hskip 1.42271pt0.9}$ $62.7_{\hskip 1.42271pt4.3}$ $85.9_{\hskip 1.42271pt3.7}$ $\textbf{80.1}_{\hskip 1.42271pt\phantom{1}1.2}$ $77.2$
CompRW $24$ $90.8_{\hskip 1.42271pt1.8}$ $80.6_{\hskip 1.42271pt2.1}$ $\textbf{76.4}_{\hskip 1.42271pt1.7}$ $\textbf{60.7}_{\hskip 1.42271pt4.4}$ $\textbf{81.6}_{\hskip 1.42271pt1.0}$ $\textbf{65.3}_{\hskip 1.42271pt3.4}$ $\textbf{88.0}_{\hskip 1.42271pt1.8}$ $79.0_{\hskip 1.42271pt\phantom{1}1.6}$ 77.8
Llama-3-8B Standard $3,4$ $91.4_{\hskip 1.42271pt1.7}$ $79.2_{\hskip 1.42271pt7.2}$ $74.0_{\hskip 1.42271pt8.0}$ $58.7_{\hskip 1.42271pt4.7}$ $76.5_{\hskip 1.42271pt2.2}$ $59.4_{\hskip 1.42271pt3.7}$ $84.0_{\hskip 1.42271pt6.6}$ $87.4_{\phantom{1}\hskip 1.42271pt5.5}$ $76.3$
Standard $12$ $\textbf{92.2}_{\hskip 1.42271pt0.6}$ $\textbf{79.6}_{\hskip 1.42271pt6.9}$ $73.1_{\hskip 1.42271pt5.0}$ $\textbf{63.3}_{\hskip 1.42271pt2.4}$ $77.5_{\hskip 1.42271pt2.2}$ $62.7_{\hskip 1.42271pt3.8}$ $87.7_{\hskip 1.42271pt2.0}$ $82.1_{\hskip 1.42271pt18.1}$ $77.3$
Calib12121212 $91.1_{\hskip 1.42271pt1.5}$ $79.2_{\hskip 1.42271pt5.8}$ $\textbf{77.9}_{\hskip 1.42271pt3.3}$ $60.5_{\hskip 1.42271pt9.3}$ $77.3_{\hskip 1.42271pt2.1}$ $65.2_{\hskip 1.42271pt3.2}$ $86.4_{\hskip 1.42271pt4.7}$ $\textbf{87.7}_{\phantom{1}\hskip 1.42271pt4.0}$ $78.2$
CompRW $12$ $90.7_{\hskip 1.42271pt2.0}$ $78.3_{\hskip 1.42271pt6.7}$ $77.2_{\hskip 1.42271pt2.9}$ $61.8_{\hskip 1.42271pt6.4}$ $\textbf{78.0}_{\hskip 1.42271pt1.8}$ $\textbf{66.9}_{\hskip 1.42271pt2.4}$ $\textbf{89.1}_{\hskip 1.42271pt1.0}$ $87.4_{\phantom{1}\hskip 1.42271pt3.8}$ 78.7
Standard $24$ $\textbf{92.2}_{\hskip 1.42271pt0.8}$ $78.2_{\hskip 1.42271pt7.2}$ $78.0_{\hskip 1.42271pt2.0}$ $\textbf{63.8}_{\hskip 1.42271pt1.8}$ $76.2_{\hskip 1.42271pt3.0}$ $66.4_{\hskip 1.42271pt2.3}$ $87.9_{\hskip 1.42271pt1.9}$ $80.6_{\hskip 1.42271pt18.9}$ $77.9$
Calib24242424 $91.7_{\hskip 1.42271pt1.4}$ $\textbf{80.0}_{\hskip 1.42271pt6.1}$ $78.3_{\hskip 1.42271pt3.4}$ $\textbf{63.8}_{\hskip 1.42271pt2.9}$ $78.1_{\hskip 1.42271pt1.7}$ $66.0_{\hskip 1.42271pt2.4}$ $86.7_{\hskip 1.42271pt4.9}$ $\textbf{87.7}_{\phantom{1}\hskip 1.42271pt3.8}$ $79.0$
CompRW $24$ $91.6_{\hskip 1.42271pt1.7}$ $79.1_{\hskip 1.42271pt6.9}$ $\textbf{78.8}_{\hskip 1.42271pt2.7}$ $63.7_{\hskip 1.42271pt3.3}$ $\textbf{78.5}_{\hskip 1.42271pt1.4}$ $\textbf{67.4}_{\hskip 1.42271pt2.6}$ $\textbf{89.5}_{\hskip 1.42271pt1.1}$ $87.4_{\phantom{1}\hskip 1.42271pt3.2}$ 79.5
Table 4: ICL accuracy of 8 classification tasks and the average accuracy (Avg.). The number after a method denotes the number of labeled data used. We run 15 prompts for each task (5 disjoint sets of $K$ labeled data and 3 templates) and report the mean accuracy and standard deviation. CompRW achieves the best average accuracy in all setups.
5.3 Results
We set $K=\{12,24\}$. Table 4 compares our component reweighting (CompRW) with standard ICL (Standard), prompt selection (PromptS), and calibration (Calib+). First, we find that simply increasing the number of demonstrations from $4$ to $24$ has limited improvements in ICL accuracy, while the longer prompt greatly increases the inference time. For example, on Llama-2-7B, Standard $24$ only improves the average accuracy by $2.6\%$ over Standard $3,4$ and the accuracy even decreases on Mistral-Instruct. Second, PromptS performs the worst in most setups, likely because it is hard to find similar examples from a small pool of $K$ examples, and a bad selection induces majority label biases. Third, both calibration (Calib+) and component reweighting (CompRW) achieve substantially better accuracy than Standard $3,4$ with little test-time overhead. Overall, CompRW achieves the best average accuracy in all setups, outperforming Standard $12$ by $6.0\%,1.8\%,2.6\%,1.4$ on Llama-2-7B, Llama-2-13B, Mistral-Instruct-7B, and Llama-3-8B, respectively, and outperforming Standard $24$ by $6.0\%,2.2\%,5.1\%,1.6\%$, respectively. We run one-tailed paired t-tests comparing CompRW with Calib+ and find that p-values $<0.05$ in all 8 setups (see Table 6), showing that CompRW performs significantly better Calib+.
6 When Do Good Components Emerge?
We study the dynamics of components during pretraining by monitoring their accuracies on 32 checkpoints of Pythia-6.9B, uniformly spaced from the first to the last checkpoint. For each checkpoint, we run 4-shot ICL on AGNews with 3 templates $\times$ 3 sets of demonstrations. The demonstrations are balanced in labels with randomly shuffled orders. Figure 3 shows the average accuracy of the 9 runs shaded by the standard deviation.
While the full model (green) fluctuates and has a large variance across prompts, the top-1 components (solid blue) achieve good accuracy at an early step and plateau quickly. We also backtrack the top-1 components of different prompts at the last checkpoint (dashed blue), monitoring how they perform on average during pretraining. We observe that they are not the top components at the early stage (there are gaps between the two blue lines before the $75k$ steps), but start to perform steadily well from the middle stage. Our findings also hold on SST2 and Pythia-1.4B (see Figure 6 in the appendix), suggesting that the model’s ability to do a task emerges before it is apparent from the full model on these tasks. [^4] On the other hand, Pythia models perform poorly on the other tasks over all checkpoints; thus, the training dynamics of the model components on challenging tasks remain unclear.
7 Related Work and Discussion
Improving ICL.
Prior work shows that ICL performance varies greatly across different choices of demonstrations and templates Zhao et al. (2021); Lu et al. (2022). Specifically, Sclar et al. (2024) and Voronov et al. (2024) find no universally better prompt template that can transfer across tasks and models, implying that it is not easy to explain ICL through prompt engineering. While several approaches, such as prompt selection Liu et al. (2022b); Chang and Jia (2023); Fu et al. (2023), prompt ensemble Min et al. (2022a); Arora et al. (2023); Voronov et al. (2024), and many-shot ICL Agarwal et al. (2024), substantially improve accuracy, they treat LLMs like black boxes without understanding the internals. Besides, they greatly increase inference time or require a large set of labeled data, which deviates from true few-shot learning Perez et al. (2021). In comparison, our paper studies this problem by looking inside the LLMs. Rather than selecting prompts, we select components in a soft, learnable way. Our method only requires $\{12,24\}$ examples and has negligible computation overhead over $4$-shot ICL at inference.
Figure 3: The ICL accuracy of the full model (green) fluctuates greatly during pretraining. However, good-performing components (T1) emerge in the early steps.
Components Interpretation.
Components interpretation studies the function of different components in a trained model Elhage et al. (2021); Shah et al. (2024), where components could be neurons Radford et al. (2017); Wang et al. (2022a); Gurnee et al. (2023), attention heads Olsson et al. (2022), and MLPs Geva et al. (2021). To analyze the components, probing Alain and Bengio (2017), knockout Geva et al. (2023); Chang et al. (2024); Li et al. (2023), patching Wang et al. (2023a); Goldowsky-Dill et al. (2023), and early decoding nostalgebraist (2020); Geva et al. (2022) are widely used techniques. For example, Li et al. (2024) train a linear probe on every attention head to discover the truthful heads inside LLMs. Michel et al. (2019) and Voita et al. (2019) prune away a large percentage of attention heads and show that only a few are critical to the performance. Hendel et al. (2023), Liu et al. (2023), Merullo et al. (2024a), and Todd et al. (2024) view ICL as compressing demonstrations into function vectors, where they remove the demonstrations and modify (patch) the LLM activations at certain layers with the function vectors at test time. Early decoding interprets the investigated components in the textual space by projecting them through the output embedding matrix Geva et al. (2022). Our model decomposition is based on early decoding and we share some similarities with prior work Yu et al. (2023); Wang et al. (2023c), especially in discovering individual components that perform well on a task. Our contributions lie in providing a new view of ICL by decomposition, which reveals the transferability of components across diverse ICL settings.
Our Method vs. Pruning.
Our method caches the direct contributions of components to the outputs through the residual stream, i.e., logits = $\sum_{j}\mathbf{g}_{j}$. Thus, removing $\mathbf{g}_{j}$, the direct contribution of the component $j$, does not alter the contributions of the other components. In comparison, pruning a component changes the activations of the other components in later layers. In A.7, we show that pruning the good-performing components identified by our method greatly hurts the accuracy, meaning that pruning also defines these components as important Michel et al. (2019).
8 Conclusion
We introduce a new perspective of ICL via decomposing the model output into the sum of individual contributions of components. We then identify three types of component characteristics across 3 LLMs and 8 classification tasks. Our extensive analyses reveal consistency in component accuracy across prompts and suggest the promising direction of improving ICL by selecting components. To this end, we propose component reweighting, which learns to scale components differently on few-shot examples. Our method achieves the best average accuracy compared to prior methods. We hope this work can deepen our grasp of LLMs while motivating more methods for practical use.
9 Limitations
Our component reweighting method requires a small set of labeled data $\mathcal{D}_{\text{train}}$ to train the component weights $w$. However, we believe it is not unreasonable to have at least $K=12$ labeled examples in total and we compare with baselines using the same $K$ examples. On the other hand, we do not compare with fine-tuning-based baselines, such as LM-BFF Gao et al. (2021), T-few Liu et al. (2022a), and LoRA Hu et al. (2021), because they usually require a larger GPU memory for training and more sophisticated early stopping criteria to prevent overfitting on few-shot examples. Another limitation is that we only experiment with classification tasks for ease of evaluation. We leave it for future work to generalize our method to generation tasks by doing decomposition and reweighting at every token during generation.
Despite similarities in model decomposition, the focus of this paper is not circuits in LLMs Wang et al. (2023a). Thus, we only have limited experiments towards mechanistic understanding of the curious components in section 3.3 and A.7. Unlike prior work that uses synthetic tasks to testify whether a head attends to certain tokens Dutta et al. (2024); Merullo et al. (2024b), we work on standard NLP benchmark datasets without obviously correct or incorrect tokens to collect answers, making mechanistic interpretation more challenging.
Acknowledgements
We thank Johnny Wei for his valuable suggestions on the paper structure. We thank Qinyuan Ye, Ryan Wang, Gustavo Lucas Carvalho, Ameya Godbole, Wang Zhu, Daniel Firebanks-Quevedo, and the anonymous reviewers for their helpful feedback. This work was funded in part by gifts from Open Philanthropy and Cisco Research, and was also supported in part by the National Science Foundation under Grant No. IIS-2403436. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.
Appendix A Appendix
A.1 Tasks and Templates
Table 11 summarizes the 13 datasets we use in the paper, where we construct balanced test sets by randomly sampling 2000 examples in each task. We form the prompts by concatenating demonstrations in a randomly shuffled order. To avoid the recency bias Zhao et al. (2021), we keep shuffling the demonstrations until the last two have different labels. For minimally conservative templates (section 4.1), Table 10 compares the contrast sets we construct on Llama-2-7B. For our case study on Task069 and Task070, we sample 3 templates from Sclar et al. (2024). Figure 7 compare the prompts of Task069 and Task070, which consist of an instruction followed by $K$ templated demonstrations. Originally, the two tasks have $\sim 4\%$ of parallel examples. To make our task transfer challenging, we discard these overlapped examples.
A.2 Label-Biased Components
We say a component is label-biased when it always predicts a certain label on the entire test set section 3.2. In this section, we focus on the most biased components in binary classification tasks, i.e., the two components that have the largest value of $(\text{logit}_{0}-\text{logit}_{1})$ and $(\text{logit}_{1}-\text{logit}_{0})$, respectively, where $\text{logit}_{0}\in\mathbb{R}$ and $\text{logit}_{1}\in\mathbb{R}$ are the LLM output logits on the two classes. We name these two components as Biased Component-0 and Biased Component-1, respectively. To understand how biased these two components are, we alter the choices of demonstrations and observe their behavior. Specifically, we consider three settings: demonstrations balanced in labels (green in Figure 5), demonstrations of all negative labels ($[0,0,0,0]$; red), and demonstrations of all positive labels ($[1,1,1,1]$; blue). We fix the template and sample 5 disjoint sets of demonstrations for each setting. Each dot in Figure 5 shows the components’ prediction on an example, and the x-axis and y-axis correspond to logit0 and logit1, respectively. A dot below the dashed diagonal line means the prediction on the example is class 0. We find that both Biased Component-0 and Biased Component-1 still insist on predicting a certain label on all examples, regardless of the labels in the prompts.
Figure 4: Transformer architecture in GPT2.
Figure 5: Each dot represents an example in the test set. The two most biased components still insist on predicting the same label on the entire test set regardless of the labels of the demonstrations.
SST2 BoolQ QQP WiC RTE MNLI AGNews ARC-Easy
Llama-2-7B $37.8$ $18.9$ $43.4$ $44.2$ $35.4$ $28.2$ $13.4$ $11.5$
Llama-2-13B $32.6$ $18.7$ $37.2$ $39.3$ $32.1$ $26.0$ $12.4$ $13.1$
Mistral-Instruct-7B $31.9$ $14.0$ $32.3$ $32.4$ $27.6$ $20.9$ $14.5$ $8.4$
Llama-3-8B $38.3$ $21.4$ $42.4$ $40.0$ $36.9$ $28.1$ $15.8$ $15.0$
Table 5: We report the average percentage of label-biased components across 15 prompts for each task. A label-biased component always predicts the same label on the entire test set.
A.3 LayerNorms
Figure 4 shows the transformer architecture in GPT2-like models. Because the layernorms inside each block are before MHA and MLP, known as Pre-LN, Eq. 1 has already taken $Ln1$ and $Ln2$ into account, and Eq. 6 only has the term for the final layernorm, LN()($\cdot$)( ⋅ ).
Both Llama-2 and Mistral model families use RMSNorm Zhang and Sennrich (2019), a layer normalization variant without centering and adding bias terms. Formally, let $x\in\mathbb{R}^{d}$ be the input, the root mean square norm LN$(x)$ is:
$\text{LN}(x)$ $=\frac{x}{\text{RMS}(x)}\odot\gamma,$ (10)
$\text{RMS}(x)$ $=\sqrt{\frac{1}{d}\sum_{i=1}^{d}x_{i}^{2}},$ (11)
where $\gamma\in\mathbb{R}^{d}$ is the affine transform parameters and $\odot$ denotes element-wise multiplication.
A.4 Tests of Significance
We run one-tailed paired t-tests to test whether CompRW outperforms Calib+ significantly. In Table 4, we have the results of 15 prompts for each task and 8 tasks in total. For each model, we aggregate the 120 accuracy scores of CompRW and Calib+, respectively, and then calculate the p-values. Table 6 shows that p-values $<0.05$ in 8/8 setups, suggesting that CompRW performs significantly better than Calib+.
Llama2-7B Llama2-13B Mistral-Ins-7B Llama3-8B
$K=12$ 0.0010 0.0002 0.0470 0.0198
$K=24$ 0.0003 0.0001 0.0027 0.0245
Table 6: The p-values $<0.05$ in all 8 setups (4 LLMs, with $K=\{12,24\}$ labeled examples), showing that CompRW performs significantly better than Calib+.
SST2 ARC-Easy
Trained Full $75.8_{\hskip 1.42271pt18.1}$ $57.5_{\hskip 1.42271pt14.4}$
Oracle-T1 $91.7_{\hskip 1.42271pt\phantom{1}0.9}$ $54.5_{\hskip 1.42271pt10.1}$
CompRW $90.8_{\hskip 1.42271pt\phantom{1}1.8}$ $79.0_{\hskip 1.42271pt\phantom{1}1.6}$
Random Full $49.7_{\hskip 1.42271pt0.7}$ $25.0_{\hskip 1.42271pt0.5}$
Oracle-T1 $55.2_{\hskip 1.42271pt0.5}$ $26.7_{\hskip 1.42271pt0.2}$
CompRW $51.4_{\hskip 1.42271pt1.7}$ $25.0_{\hskip 1.42271pt0.7}$
Table 7: Comparing the ICL accuracy between pretrained (Up) and randomly initialized (Down) Llama-2-7B. The top-1 component (Oracle-T1) and CompRW perform near random on the untrained model.
A.5 Do Good-Performing Components Exist in Randomly Initialized Models?
Ramanujan et al. (2020) find that untrained subnetworks can perform on par with a ResNet-34 trained on ImageNet. Similarly, Zhang and Bowman (2018); Hewitt and Liang (2019) show that representations of randomly initialized language models yield a strong baseline for probing tasks. In this section, we investigate (1) whether good-performing components still exist in a randomly initialized LLM, and (2) how CompRW method performs using component activations extracted from the randomly initialized LLM.
We run 4-shot ICL with 15 prompts and report the average accuracy and standard deviation. For CompRW, we use the same 4 demonstrations and 20 more examples for reweighting. Table 7 shows that the best-performing component (Oracle-T1) in a randomly initialized Llama-2-7B still performs poorly on SST2 and ARC-Easy. While CompRW has substantial improvement over Full on the pretrained model, it has no effect on the randomly initialized model. We conclude that good-performing components do not exist in a randomly initialized LLM and our CompRW method relies on the pretrained component activations to perform well.
A.6 More Results on Transferability
In section 4, we study the transferability of components across different choices of demonstrations and templates. Here, Table 9 shows the full results on all LLMs and tasks. We observe the same findings as Table 2: component accuracies agree well across randomly sampled demonstrations, but have much weaker agreements across randomly sampled templates. Because constructing minimally-contrastive templates requires non-trivial manual efforts, we only build contrast sets for 5 tasks on Llama-2-7B (shown in Table 2), where these tasks have the largest variances across templates.
A.7 Pruning Good and Bad Components
Our method studies a component using its cached direct contribution to the output, whereas Michel et al. (2019) (pruning) zeroes out the activations of a component in the forward pass and thus indirectly changes the activations of other components in the upper layers. They consider a component important if pruning it causes large drops in task performance. In this section, we investigate the intersection between our method and pruning.
[…]
SST2 RTE MNLI AGNews
Full Model $\textbf{75.8}_{\hskip 1.42271pt18.1}$ $\textbf{68.9}_{\hskip 1.42271pt3.2}$ $34.4_{\hskip 1.42271pt1.7}$ $70.0_{\hskip 1.42271pt19.9}$
Prune-T50 ${\color[rgb]{1,0,0}53.4}_{\phantom{1}\hskip 1.42271pt8.8}$ ${\color[rgb]{1,0,0}57.8}_{\hskip 1.42271pt6.4}$ $34.3_{\hskip 1.42271pt3.2}$ ${\color[rgb]{1,0,0}26.8}_{\phantom{1}\hskip 1.42271pt2.5}$
Prune-B50 $72.3_{\hskip 1.42271pt14.8}$ $65.4_{\hskip 1.42271pt5.8}$ $\textbf{35.7}_{\hskip 1.42271pt4.0}$ $\textbf{72.6}_{\hskip 1.42271pt15.7}$
Table 8: Comparing the accuracies of the full Llama-2-7B model and pruning the top/bottom 50 components. We run 15 prompts for each task and report the average accuracy and standard deviation. We color the numbers red when there is a large drop in accuracy.
A.8 Training Details and Hyperparameters
For both CompRW and Calib+  methods, we train a linear layer on $\mathcal{D}_{\text{train}}$ with stochastic gradient descent. Because we do not have an additional dev set to tune the hyperparameters, we use the same hyperparameters on all the tasks and models and do early stopping based on the loss and accuracy on $\mathcal{D}_{\text{train}}$. Specifically, we set learning rate $=0.05$ for both methods and $\lambda=0.1$ for the L1 regularization term in CompRW. We run all our ICL experiments on a single RTX A6000 GPU (48G). Both the component reweighting and calibration training processes can be run on a single i7 CPU within a minute.
A.9 Models
We use the model checkpoints on Hugging Face, meta-llama/Llama-2-7b-hf, Llama-2-13b-hf, mistralai/Mistral-7B-Instruct-v0.1, and meta-llama/Meta-Llama-3-8B.
SST2 BoolQ QQP WiC RTE MNLI AGNews ARC
Correlation Llama-2-7B
(1) Demo 0.81 0.84 0.60 0.65 0.75 0.65 0.89 0.88
(2) Temp 0.40 0.16 0.03 0.15 0.19 0.09 0.68 0.44
IoU
(1) Demo 0.36 0.74 0.27 0.21 0.53 0.24 0.63 0.70
(2) Temp 0.12 0.01 0.01 0.03 0.05 0.01 0.20 0.20
Correlation Llama-2-13B
(1) Demo 0.83 0.84 0.63 0.67 0.78 0.73 0.91 0.91
(2) Temp 0.57 0.30 0.09 0.19 0.28 0.16 0.76 0.55
IoU
(1) Demo 0.26 0.71 0.31 0.18 0.46 0.39 0.55 0.65
(2) Temp 0.21 0.11 0.07 0.01 0.21 0.07 0.25 0.30
Correlation Mistral-Instruct-7B
(1) Demo 0.88 0.91 0.72 0.75 0.87 0.82 0.92 0.97
(2) Temp 0.58 0.44 0.19 0.26 0.40 0.30 0.77 0.60
IoU
(1) Demo 0.39 0.59 0.27 0.29 0.50 0.45 0.68 0.80
(2) Temp 0.10 0.17 0.06 0.05 0.17 0.09 0.29 0.22
Correlation Llama-3-8B
(1) Demo 0.85 0.88 0.70 0.73 0.80 0.81 0.89 0.95
(2) Temp 0.55 0.39 0.26 0.25 0.31 0.23 0.67 0.52
IoU
(1) Demo 0.42 0.56 0.28 0.25 0.46 0.52 0.65 0.68
(2) Temp 0.15 0.12 0.09 0.07 0.08 0.05 0.34 0.27
Table 9: Full results of the average correlation and IoU between (1) two random sets of demonstrations and (2) two randomly sampled templates.
Task Templates Labels Accuracy
SST-2
[…]
negative/positive $50.6\pm 0.7$
[…]
negative/positive $72.7\pm 6.1$
BoolQ
[…]
No/Yes $52.5\pm 2.0$
T2 Based on the following passage, {question}? {passage} Answer: {label}
No/Yes $66.7\pm 2.1$
QQP
T1 Are the questions "{sent1}" and "{sent2}" asking the same thing? {label}
no/yes $54.3\pm 1.1$
T2 Are the questions "{sent1}" and "{sent2}" asking the same thing? {label}
No/Yes $68.7\pm 4.1$
AGNews
[…]
World/Sports/Business/Technology $43.9\pm 8.7$
T2 {text} Is this a piece of news regarding World, Sports, Business, or Technology? {label}
World/Sports/Business/Technology $88.5\pm 0.8$
Table 10: We construct minimally contrastive templates that only differ slightly (colored in red) but yield large differences in 4-shot ICL accuracy on Llama-2-7B. We report the average accuracy and standard deviation across 5 ICL runs with different demonstrations under the same template.
Dataset Task # Classes
SST-2 Socher et al. (2013) Sentiment Analysis 2
Yelp-polarity Zhang et al. (2015) Sentiment Analysis 2
BoolQ Clark et al. (2019) Yes/No QA 2
BoolQ Contrast Set Gardner et al. (2020) Yes/No QA 2
QQP Wang et al. (2018) Paraphrase Identification 2
WiC Pilehvar and Camacho-Collados (2019) Word Sense Disambiguation 2
RTE Wang et al. (2018) Natural Language Inference 2
MNLI Williams et al. (2018) Natural Language Inference 3
MedNLI Romanov and Shivade (2018) NLI in Medical Domain 3
AGNews Zhang et al. (2015) Topic Classification 4
ARC-Easy Clark et al. (2018) Multiple-Choice QA 4
Task069 Mishra et al. (2022); Wang et al. (2022b) Abductive NLI 2
Task070 Mishra et al. (2022); Wang et al. (2022b) Abductive NLI 2
Table 11: Summary of all the datasets.
Figure 6: 4-shot ICL accuracy on different pretraining checkpoints. We compare the full model (green) with the top-1 (solid blue) and bottom-1 (red) components. The dashed blue line tracks how the top-1 components of the last checkpoint (Last-T1) perform across time.
Figure 7: Comparing the prompts of Task069 and Task070. We apply the templates of Sclar et al. (2024) and prepend the task instructions before $K$ demonstrations. We ensure that the two tasks do not have parallel examples to make the transfer experiment (section 4.3) challenging.
Figure 8: Each dot represents a component (attention head: blue; MLP: orange) under 4-shot ICL on Mistral-Instruct-7B. The x-axis shows how often a component predicts label 1 across the test data of a binary task.